please help!!

How to build a machine which has the sole purpose of turning on a light bulb?

minimum requirement:
8 steps
1 light bulb
size shouldnt be big


machine type can be anything. it doesnt matter if its run by electric circuits, magnetic field or something else.
Answer #1

In today's time when cost of electricity is very high so lets light a bulb using solar energy ,here in steps i will show how to do that

Suppose you have to light a bulb of 18 watt ac at the place banglore in India (you can choose your own place of living)

Design steps

Step01:-identify the load requirement :-

18 watts ac load for 10 hr backup :

Hence Watt hr requirement

Wh=18*10=180wh

Step02:- Choosing number of back up days for battery:-

11 year data of place suggests to take 3 days backup for the place for 12 month non stop operation and 3 days battery 4.239 kwh/m2/day (Get your place latitude and longitude and get the same value of minimum daily energy for your place ).

Battery size required in (Ah)

Ah=wh*na/Vnom

Ah=180*3/12

Ah=45

Here ,

Vnom =Nominal rated voltage of system.

na =Number of backup days for system designed.

wh =Watt hour of load required.

Means we need battery which can have ratings of 12 volt and 15 ah of ampere hours.

For safety and market availability i have chosen (C-10, 42 ah ,12 volt) battery.Type of battery chosen is Exide lead acid sealed battery .

Although one can prefer solar tubular battery with higher ah in same cost but maintenance issue will be there due to water issue in those battery ...to avoid that i have chosen sealed battery in this experimental setup.

Step03:-Designing solar charge controller

The role of controller is to make sure adequate amount of charge is transfer to batter, load from PV panel.

Charge controller chosen is 10 amp 12 volt to maintain the charge between pv panel load and for load current of approx 4 amp peak .

The Icc(charge controller current),

Icc=4*1.25

Icc=5 amp.

here, 25 percent safety for losses and microcontroller ckt plus cable losses is taken into consideration.

So for market availability one can choose 5 amp 12 volt for dc application only but 12 volt 10 amp is better in case of overloading or malfunction of some component.

  

Step 04:-Choosing panel size :-

Now for connection purpose i have bought 10 meter 4mm2 dc cable @ market price of 1100 rupees. ,one 200 watt load serving inverter and panel.Now say inverter gives me 85 percent efficiency and overall cable losses is say makes overall efficiency of system to be 75 percent for safer side.

Hence panel must surve load hr of 540wh/0.75=720wh/day load.

Now say we take two  100 watt PV panel in series with

Short circuit current

           

Isc=6.2 max ,hence for that charging current that panel will give is

Icharging=0.9*0.85*6.2*2

Icharging=4.743 amp

HEre above 0.9,0.85 are efficency and 2 is number of panels

Now for the banglore place (Choose your place value here) for the period of april and may i have Hat _min=4.238 kwh/m2/day Hence for 1 kwh standard insolation this is equivalent to 4.238 hours minimum sun hours .

Implies that minimum ampere hours given by panel is

Ahmin=4.2338*Icc

Ahmin=5.635*4.743*2

Ahmin=26.7975*2 ah=53.45 ah what panel will give

Recall from above i needed 720 wh daily

For 12 volt system i get720/12=60Ah watt hour of

See 53.45 ah is what we get from panel daily while 60 ah we need hence we can design the system for the above approximation with 7 ah approx less on daily basis hence we will get approx 9 to 10 hour of light at night from battery choosen for 18 watt ac bulb to light up at all weather condition .

//If you donnt understand anything let me know in comment section

// i will try and help